Simplify the following boolean algebra equation only with the xor logic gate!

Please answer clearly

a. (AB+CD)(A+B)(AB)+(C+B)(C+B)+(A+1)A(B+C+D)AB(A’+B) b. (((CD+CD)+((A’B)+(AB))*(((A’B)+(AB))+ (CD’+CD)))
Answer #1

@ A.B = Ãtà AB + C D (Ā+B)(AB) + (C+B) (C+B) +(A+1) A (B+C+D) 0 @ AB(A+B) Demorgans law AB=Ā+B Ā.B = A+B A+BC = (A+B) (A+C® ((CD + CD + (ñ8+AZ)( 26+) +[C0+))) A + B = AB CD+CD = CAD ĀB+AB AB. AB Ā B + A B. AŘ + - * (@O)1) (1 + (HD) CODTD+CÒ COD= T

In given question b there is extra bracket

If I was wrong anywhere please comment so I can improve answer